The McGraw Company is accumulating data to be used in preparing its annual profit plan for the coming year. The cost behavior pattern of the maintenance costs must be determined. The accounting staff has suggested that linear regression be employed to derive an equation in the form of y = a + bx for maintenance costs. Data regarding the maintenance hours and costs for last year and the results of the regression analysis are as follows: (CMA adapted)

Hours of Activity Maintenance Costs
January 480 $ 4,200
February 320 3,000
March 400 3,600
April 300 2,820
May 500 4,350
June 310 2,960
July 320 3,030
August 520 4,470
September 490 4,260
October 470 4,050
November 350 3,300
December 340 3,160
Sum 4,800 $ 43,200
Average 400 3,600
A coefficient 684.65
B coefficient 7.2884
Standard error of the a coefficient 49.515
Standard error of the b coefficient 0.12126
Standard error of the estimate 34.469
R2 0.99724
T-value a 13.827
T-value b 60.105


Using the high-low method to estimate cost behavior, 420 maintenance hours in a month would mean the maintenance costs would be budgeted at:

  • $3,600.

  • $3,150.

  • $3,780.

  • $3,720.
Answer #1

Variable cost per unit = (Highest activity cost - Lowest activity cost)/(Highest activity - Lowest activity)

= (4,470-2,820)/(520-300)

= 1,650/220

= 7.5 per hour

Fixed cost = Highest activity cost - Highest activity x Variable cost per hour

= 4,470 - 520 x 7.5

= 4,470 - 3,900

= $570

estimate cost behavior, 420 maintenance hours in a month would mean the maintenance costs would be budgeted at=

= Fixed cost + Variable cost per hour x Activity hour

= 570 + 7.5 x 420

= 570+3,150

= $3,720

Fourth option is correct option.

Kindly give a positive rating if you are satisfied with this solution and please ask if you have any query.

Thanks